OpenStudy (anonymous):

The Senate passed a bill by a ratio of 8 to 1. If 90 senators voted in all, how many voted for the bill?

OpenStudy (anonymous):

Choices: 80 89 11.25 10

OpenStudy (anonymous):

I think it is 89

OpenStudy (anonymous):

no it is not that

OpenStudy (anonymous):

oh, ok...ill go over it again

OpenStudy (anonymous):

8 to 1 means for every nine that voted in favor, 8 voted for it and 1 against

OpenStudy (anonymous):

8/9 * 90 = 80

OpenStudy (anonymous):

8 to 1 gives 8/(8+1) = x/90 and then just solve for x.

OpenStudy (anonymous):

what i mean to say is that for every 9 that VOTED , 8 voted for and 1 voted against

OpenStudy (anonymous):

so by the same ratio, if 90 voted, then 80 voted for and 10 voted against

OpenStudy (anonymous):

oh ok @satellite73

OpenStudy (anonymous):

8/(8+1) = x/90 simplifies to 8/9 = x/90

OpenStudy (anonymous):

ok!!! :)

OpenStudy (anonymous):

thank you so much! :) <3

OpenStudy (anonymous):

and if 45 voted, then 40 voted for it and 5 against etc

OpenStudy (anonymous):

which then gives x = (8/9)(90)

OpenStudy (anonymous):

You're welcome!